What is the moment of inertia of the object about an axis at the center of...

momentofinertia2new

What is the moment of inertia of the object about an axis at the center of mass of the object? (Note: the center of mass can be calculated to be located at a point halfway between the center of the sphere and the left edge of the sphere.)
Answer #1

mass of the sphere ms

mass of the rod mr

radius of the sphere R

length of the rod L.

Moment of inertia of the sphere about the center of mass of the system (by parallel axis theorem)

I_s=\frac{2}{5}m_s R^2+m_s L^2

I_s=\frac{2}{5}m_s R^2+m_s \left (\frac{R}{2} \right ) ^2

I_s=\frac{13}{20}m_s.R^2

Moment of inertia of rod about the center of mass of the system (by parallel axis theorem)

I_r=\frac{1}{12}m_r L^2+m_r*\left ( \frac{L}{2}+\frac{R}{2} \right )^2

I=I_r+I_s=\frac{13}{20}m_s.R^2+\frac{1}{12}m_r L^2+m_r*\left ( \frac{L}{2}+\frac{R}{2} \right )^2

just substituing values to find the solutution.